# From the light curves of Long Period Variables to their evolution along the AGB

### v1

@@ Abstract
A sample of 116 Long Period Variables is studied with the aim to reveal relations between the properties of their light curves and of their evolution along the Asymptotic Giant Branch (AGB). Each light curve is carefully scrutinized and its properties are summarized in a set of five parameters. The picture that had been sketched in an earlier work is clarified and new features are revealed. In particular, the evolution on the AGB of stars having not yet experienced a third dredge-up event is shown to go together with a closely related path followed by the light curve in the parameter space. Moreover, evidence is found for the existence of two different types of light curves for stars having experienced dredge-up events strong enough for their impact to be detectable. Interpretations are proposed, however often too speculative to be firmly ascertained: the results presented raise several unanswered questions and point to the need for new observations and analyses, suggesting that light curves still carry more messages and information than has been currently possible to decrypt.
@@
show]mainhattanhdpt@gmail.com
@@ 1 Introduction
Fifty years ago, M. Schwarzschild (1975) argued that convection was probably a major contributor to the mass loss rate of Red Giants and was likely to materialize in the form of a very few convective cells, each covering typically a steradian. His argument was based on a scaling of the Sun’s supergranulation. Today’s understanding of solar supergranulation has somewhat progressed (F. Rincon & M. Rieutord, 2018), but the argument still applies in broad terms and observations using VLTI and ALMA, together with VLBI detections of masers, have confirmed the existence of such convective cells. In broad terms, it seems well established that the mass loss of oxygen-rich stars has its source in radial movements associated with pulsations and the shock waves which they generate, both at the global star and convective cell levels, whether at steradian scale with life times of years or at smaller scales with much shorter life times (C. Paladini et al., 2018; A. Rosales-Guzmán et al., 2024; W. Vlemmings et al., 2017; W. H. T. Vlemmings et al., 2019; W. Vlemmings et al., 2024; P. Darriulat et al., 2024).
@@ 2 The sample of light curves and their parameterization
A particularly useful parameter largely used in Paper I is the asymmetry between the durations of the ascending and descending branches. In the case of sample B we preferred to use the equivalent parameter . Depending on the method used to measure such asymmetry, different systematic uncertainties are attached to and . In the present article we systematically use the former. We recall that in Paper I, we noted important differences between the properties of the light curves of spectral type Mno according to the value of : we defined a first Mno family, Mno1, as having . In contrast with curves of the second family, Mno2, their light curves were found to display no hump on the ascending branch, to be more symmetric with between 0.50 and 0.55 (compared with between 0.57 and 0.65 for Mno2), to have smaller periods, below 300 days, to be more regular and to have smaller colour indices than Mno2. They sit on the left-hand side of the colour vs period diagrams first produced by S. Uttenthaler et al. (2019), while Mno2 occupies also the right-hand side. When observed in infrared, the oscillation amplitude is less than 20% of the value in the visible, while it exceeds 25% for Mno2.
@@
A well-known feature displayed by the light curves is a clear difference between profiles having a narrow maximum and a broad minimum and profiles having a broad maximum and a narrow minimum. The left and central panels of Figure 1 display two extreme cases. Such distinction had already been made by T. Lebzelter (2011), however without producing very significant results. We searched for parameters that quantify reliably this feature and found two good candidates, which we call and (Figure 1, right); is the width of the curve, relative to the period, at an apparent luminosity equal to 20% of the luminosity amplitude of the oscillation; is the ratio between the luminosity measured from minimum light and integrated over a period and the product of its value at maximum light by the period. Pulsations with relatively short values of the parameters mean a star that is mostly dim during its cycle, pulsations with relatively large values mean a star that is mostly bright during its cycle. We find that, in general, the two parameters are closely related, with, on average, 1.5 . Apart from a small difference between oxygen-rich and carbon-rich stars, we did not find significant cases of deviation between the information provided by the two parameters and considered it sufficient, for the purpose of the present study, to use a single width parameter , defined as the geometric mean between and , . The main difference between parameter and similar parameters commonly used in the published literature, including Paper I, is its being evaluated from the luminosity rather than from the magnitude. This has a trivial but strong effect when the oscillation amplitude is large; a good illustration is the case of chi Cyg which has a very low value of , 0.18, while its profile, studied in Paper I, is broad and displays a hump on the ascending branch.
@@ 3.1 The Mno spectral type
In the present section we study the light curves of stars having an Mno spectral type, as listed in Table 1. More precisely, we exclude four such stars that display peculiarities that will be discussed in Subsection 4.2 (W Dra, W Hya, T Ari and U CMi), and we include a star of spectral type M2, RT Cyg, for which technetium had been searched for and found unlikely to be present. Figure 2 illustrates the distributions of the curves in the parameter space by displaying projections on the vs , vs , vs , vs and vs planes. It reveals a remarkable dependence on the temperature of the star, as measured by the value of index for an M- spectral type. As increases, namely as the star cools down, decreases and all other parameters increase. This evolution, for which Table 2 gives evidence, is particularly clear in the cases of , , and , slightly less in the case of and unclear in the case of .
@@
When the star has a large enough initial mass, it will at some point experience sufficiently strong TDU for its spectral type to change to Myes and eventually S. If the above interpretation of the evolution of its light curve is correct, we should therefore expect the associated light curves to populate the low part of the parameter space, say below 0.3. In the next subsection, we use this argument as guideline to seek an answer to the second question that we need to address: what will the stars in our Mno sample evolve to?
@@
The study presented in the preceding subsection suggested that the stars of Mno spectral type listed in Table 1, when initially sufficiently massive, would evolve to the Myes spectral type with light curves of small width parameter, say 0.3. Table 3 lists all curves of our sample, excluding those listed in Table 1, that have an M or S spectral type and for which the light curve has a width parameter 0.3. Their distribution in the parameter space is displayed in Figure 3. As decreases, the period, the oscillation amplitude and the colour index tend to increase. Remarkably, a dozen of curves display large irregularity parameters, well in excess of 0.6, in strong contrast with the Mno sample where a single curve, that of R Cas, has a large value of ; it has also extreme values of , and , suggesting that it is close to evolve to an Myes spectral type. The presence of a large group of irregular curves in the present sample suggests that some are evolving toward a state of unstable pulsation and might be leaving the domain of stable pulsation before becoming carbon-rich stars. Indeed, F. L. Schöier & H. Olofsson (2001), in a study of a large sample of bright carbon stars, have shown that only a quarter of them pulsate as Mira variables.
@@
The curves selected in the preceding subsection were required to have 0.3. In order to see whether such a criterion is justified, we need to study the curves of the oxygen-rich stars having 0.3, excluding of course the curves of Mno type listed in Table 1, and see how different they are from those listed in Table 3. They are listed in Table 4 and Figure 4 displays their distribution in the parameter space. At first glance, Figure 4 seems to suggest that the 0.3 curves are successors of the Mnoi-j curves having 4, while the 0.3 curves are successors of the Mnoi-j curves having 3; moreover, the location of the 0.3 curves between the Mno and carbon-rich samples may suggest that the stars having 0.3 curves will evolve to the carbon-rich variables of the present sample. A close inspection of the properties of the curves listed in Table 4 is therefore necessary to understand whether such crude impressions are or not justified.
@@
The trend toward low colour indices is already apparent in the Table 4 sample of the twelve curves having 0.4. Indeed, out of the 548 curves selected by P. Merchán Benítez et al. (2023), 40 have [3.4]-[22]1.25, of which 19 are of S spectral type, 18 of M spectral type and 3 of C spectral type. Of these, 10 are part of the Table 4 sample and only 5, all of S spectral type with between 0.24 and 0.27, are part of the much larger Table 3 sample. This is evidence for a group of stars of very low colour indices, mostly of S, but also Myes spectral types, having curves of large parameter, to own a specific identity that must be associated with well-defined physical properties. A likely interpretation is offered by O. Hashimoto et al. (1998) in the case of R Hya, a star of the Table 4 sample having an M6- 9 spectral type, a colour index 0.9 and a broad curve profile, 0.36. The claim is that mass-loss has recently nearly stopped, causing the circumstellar envelope to expand away from the star and become too cold for the silicate dust grains to significantly emit in the infrared. Recent observations and analyses (L. Decin et al., 2008; A. Baudry et al., 2023; M. Joyce et al., 2024; R. Zhao-Geisler et al., 2012) have confirmed the validity of this interpretation and suggested that the cause of the absence of significant mass-loss is the likely occurrence of a thermal pulse some 200 years ago. Models such as by Y. A. Fadeyev (2024) quote a large stellar mass of 4 to 5 solar masses.
@@
In summary, an interpretation of the Table 4 curves as associated with large initial mass stars near the end of the TP-AGB, having light curves evolving from early Mno type (4) to the carbon-rich Table 5 sample is plausible but far from being firmly established and raises several questions that deserve careful consideration.
@@ 3.4 Carbon-rich stars
The light curves of the carbon-rich stars of the present sample are listed in Table 5 and their distribution in the parameter space is illustrated in Figure 4. The width parameter exceeds 0.3. A clear correlation is seen between and and a weaker one between and and and . The analyses presented in the preceding subsections have left open a reliable identification of their progenitors. They found plausible that it hosts members of the Table 4 sample. But they could not exclude that it would also host members of the Table 3 sample. In particular, transition from Table 3 to Table 5 might occur via a regime of unstable pulsation: such curves have lost memory of the region where they come from: both =0 and =1 curves are curves of constant luminosity.
@@
J. Bergeat & L. Chevallier (2005) have collected data related to a large sample of carbon-rich stars, of which 15 are part of the Table 5 sample. Figure 5 displays the dependence of the parameters and of their light curves on effective temperature and mass loss rate . If one excludes four low curves that are close to the domain covered by the Table 4 sample, V CrB, U Cyg, RZ Peg and R Ori, tends to increase with temperature and decrease with mass loss rate while displays the opposite trend. Such a behaviour would therefore suggest that the stars having large light curves evolve in the direction of decreasing when becoming cooler. This trend is confirmed by the dependence of on the spectral index for spectral types C: for 6, namely warmer stars, 0.530.07 while for 6, namely for cooler stars, 0.390.04. This is however excluding R For, a clear outlier having a C4 spectral type and 0.35.
@@
To the extent that we can take as granted that carbon stars keep cooling down (P. Marigo, 2002), this suggests that the stars of the Table 5 sample having curves of large value enter it from the large side and evolve toward smaller values of . If such is the case, they would probably have a Table 3 origin: the larger , the smaller the oscillation amplitude (Figure 4), namely the closer from a regime of unstable pulsation. Figure 6 illustrates what such a transition would look like by displaying projections on the vs , vs , vs and vs planes of the distributions in parameter space of the Table 3 and Table 5 samples. It shows that, apart from and , a transition from Table 3 to Table 5 is reasonably continuous in the other parameters. The discontinuity in amplitude is not surprising: absorption of the dust envelope is known to be larger for carbon-rich rather than oxygen-rich stars.
@@
RS Cyg and V Oph, both parts of the Figure 5 sample, have large values of , 0.66 and 0.52, respectively. They would then be likely successors of the Table 3 sample. RS Cyg has a high temperature of 3100 K and a low mass loss rate of 210-7 solar masses per year (R. Guandalini et al., 2006). Its light curve has been extensively studied and commented by R. R. Cadmus (2022, 2024). It often displays a hump on the ascending branch, close to maximum light, possibly producing a double-maximum profile. V Oph has been the target of detailed VLT observations in the infrared (K. Ohnaka et al., 2007; G. Rau et al., 2019; J. Hulberg et al., 2025) suggesting an effective temperature of 2700 K while Figure 5 quotes a temperature of 3010 K and a mass loss rate of 1.410-7 M⊙yr-1. However, the identification of the origin of curves having lower values of turns out to be difficult. BH Cru and LX Cyg were shown by S. Uttenthaler et al. (2011, 2016) to have recently switched from spectral type S to C: one would naively expect to find them near an edge of the domain covered by the carbon-rich sample in the parameter space; but such is not the case, their curves have values of (0.47 and 0.46, respectively) that sit in the middle of those of the Table 5 sample, a bit on the high side: they may be successors of the Table 3 sample as well as of the Table 4 sample, although the value of the temperature (3000 K, A. A. Zijlstra et al., 2004) rather favours the former.
@@
S Cep, with 0.40, has an effective temperature of 2100 K (W. Nowotny et al., 2005, 2010), leaving open both a Table 3 and a Table 4 origin. Such is indeed the case for all curves having values of this order or smaller. The 13 curves having 0.4 cover rather uniformly the parameter space, failing to reveal a clear separation between two groups of curves. These examples illustrate the difficulty of producing reliable interpretations of the evolution of the stars having curves in the Table 5 sample: the likely coexistence of Table 3 and Table 4 successors is a major complication and a reliable distinction between them is difficult.
@@
Finally, we recall that we have shown in Paper I the dependence of and on and remarked that carbon-rich stars and oxygen-rich stars follow similar trends, the former having slightly lower values of and of than the latter for a same value of . Figure 7 illustrates this issue in the present sample, displaying separately the samples of Tables 1, 3, 4 and 5. The Table 1 distributions illustrate the transition from what was defined as first and second families of Mno curves in Paper I, with both and increasing; the location of the curve of R Cas (0.84) confirms the suggestion that was made in Subsection 3.2 that the star is in the process of switching to the Myes spectral type. The same trend is observed in the distributions of the Table 3 sample, with a global shift toward larger and ; a clear shift toward large values of gives evidence for the increased instability of the pulsation regime. The concentration in small values of both and illustrates the difference between the curves of the Table 4 sample and the others. The distributions displayed by the Table 5 sample confirm the comments made in Paper I.

### v2

@@ Abstract
A sample of 116 Long Period Variables is studied with the aim to reveal relations between the properties of their light curves and of their evolution along the Asymptotic Giant Branch (AGB). Each light curve is carefully scrutinized and its properties are summarized in a set of five parameters. One of these, which measures the relative width of the light maxima and minima, is found to be particularly efficient at revealing significant correlations with the current state of the star on the AGB. The picture that had been sketched in an earlier work is clarified and new features are revealed. In particular, the evolution on the AGB of stars having not yet experienced a third dredge-up event is shown to go together with a closely related path followed by the light curve in the parameter space. Moreover, evidence is found for the existence of two different types of light curves for stars having experienced dredge-up events strong enough for their impact to be detectable. For some, probably associated with higher initial mass stars, the light maximum becomes broader as the star evolves along the AGB, while for others it becomes narrower. Interpretations are proposed, however often too speculative to be firmly ascertained: the results presented raise several unanswered questions and point to the need for new observations and analyses, suggesting that light curves still carry more messages and information than has been currently possible to decrypt.
@@
show]mainhattandhpt@gmail.com
@@ 1 Introduction
Fifty years ago, M. Schwarzschild (1975) argued that convection was probably a major contributor to the mass loss rate of Red Giants and was likely to materialize in the form of a very few convective cells, each covering typically a steradian. His argument was based on a scaling of the Sun’s supergranulation. Today’s understanding of solar supergranulation has somewhat progressed (F. Rincon & M. Rieutord, 2018), but the argument still applies in broad terms and observations using VLTI and ALMA, together with VLBI detections of masers, have confirmed the existence of such convective cells. In broad terms, it seems well established that the mass loss of oxygen-rich stars has its source in radial movements associated with pulsations and the shock waves which they generate, both at the global star and convective cell levels, whether at steradian scale with life times of years or at smaller scales with much shorter life times (C. Paladini et al., 2018; A. Rosales-Guzmán et al., 2024; W. Vlemmings et al., 2017, 2019, 2024; P. Darriulat et al., 2024).
@@ 2 The sample of light curves and their parameterization
A particularly useful parameter largely used in Paper I is the asymmetry between the durations of the ascending and descending branches. In the case of sample B we preferred to use the equivalent parameter . Depending on the method used to measure such asymmetry, different systematic uncertainties are attached to and . In the present article we systematically use the former. We recall that in Paper I (Sections 4.1 and 7.1), we noted important differences between the properties of the light curves of spectral type Mno according to the value of : we defined a first Mno family, Mno1, as having . In contrast with curves of the second family, Mno2, their light curves were found to display no hump on the ascending branch, to be more symmetric with between 0.50 and 0.55 (compared with between 0.57 and 0.65 for Mno2), to have smaller periods, below 300 days, to be more regular and to have smaller colour indices than Mno2. They sit on the left-hand side of the colour vs period diagrams first produced by S. Uttenthaler et al. (2019), while Mno2 occupies also the right-hand side. When observed in infrared, the oscillation amplitude is less than 20% of the value in the visible, while it exceeds 25% for Mno2.
@@
A well-known feature displayed by the light curves is a clear difference between profiles having a narrow maximum and a broad minimum and profiles having a broad maximum and a narrow minimum. The left and central panels of Figure 1 display two extreme cases. Such distinction had already been made by T. Lebzelter (2011), however without producing very significant results. We searched for parameters that quantify reliably this feature and found two good candidates, which we call and (Figure 1, right); is the width of the curve, relative to the period, at an apparent luminosity equal to 20% of the luminosity amplitude of the oscillation; is the ratio between the luminosity measured from minimum light and integrated over a period and the product of its value at maximum light by the period. Pulsations with relatively short values of the parameters mean a star that is mostly dim during its cycle, pulsations with relatively large values mean a star that is mostly bright during its cycle. We find that, in general, the two parameters are closely related, with, on average, 1.5 . Apart from a small difference between oxygen-rich and carbon-rich stars, we did not find significant cases of deviation between the information provided by the two parameters and considered it sufficient, for the purpose of the present study, to use a single width parameter that accounts for the sensitivity carried by each of and . We define as the geometric mean of and , . The main difference between parameter and similar parameters commonly used in the published literature, including Paper I, is its being evaluated from the luminosity rather than from the magnitude. This has a trivial but strong effect when the oscillation amplitude is large; a good illustration is the case of Cyg which has a very low value of , 0.18, while its profile, studied in Paper I, is broad and displays a hump on the ascending branch.
The spectral type, Mi-j, S or C, is a reliable tracer of the evolution of the star along the AGB. The transition from M to S and from S to C, when it occurs, reveals unambiguously the last steps of such evolution before leaving the AGB. It is associated with a drastic change of the nature of the outer atmosphere, from oxygen rich to carbon-rich, resulting in a dramatic modification of its dust content. An essential milestone in the M spectral type is the presence or absence of technetium that reveals reliably the transition from the E-AGB to an advanced state of TP-AGB experiencing strong enough TDU events. Unfortunately, only few stars of M spectral type have been searched for technetium and the difficulty of the observation often prevents a clear answer to be obtained.
@@ 3.1 The Mno spectral type
In the present section we study the light curves of stars having an Mno spectral type, as listed in Table 1. More precisely, we exclude four such stars that display peculiarities that will be discussed in Subsection 4.2 (W Dra, W Hya, T Ari and U CMi), and we include a star of spectral type M2, RT Cyg, for which technetium had been searched for and found unlikely to be present. Figure 2 illustrates the distributions of the curves in the parameter space by displaying projections on the vs , vs , vs , vs and vs planes. It reveals a remarkable dependence on the temperature of the star, as measured by the value of index for an Mi-j spectral type. As increases, namely as the star cools down, decreases and all other parameters increase. This evolution, for which Table 2 gives evidence, is particularly clear in the cases of , , and , slightly less in the case of and unclear in the case of .
@@
When the star has a large enough initial mass, it will at some point experience sufficiently strong TDU for its spectral type to change to Myes and eventually S. To the extent that TDUs modify only progressively its outer atmosphere, we can expect the light curve of such a star to keep evolving in the direction of decreasing values of . Yet, we see that while some curves of Myes and S type stars have indeed small enough values to be consistent with such a scenario, others have values of that are clearly too large, exceeding up to 0.4 in the case of R Cam and S UMa. This is surprising and requires an explanation. One may think of several reasons for the simple scenario to fail in such cases: the star may have left the Mno spectral type very early and its curve did not have time to evolve toward lower values but evolved instead toward larger values; or the curve of the star, when it had an Mno spectral type, was too irregular to be part of our sample; or it simply does not evolve as progressively as we assume… In any case, in order to clarify this important issue, it is convenient to discuss separately curves that may evolve toward smaller values of and curves that seem instead to evolve toward larger values of . To this end, we define a value of 0.3 to make the distinction between the two cases; this is of course an arbitrary choice, aimed only at easing the discussion, and should not be misinterpreted as a physically meaningful number.
@@
In line with the discussion presented in the preceding sub-section, we start with the study of the curves having 0.3 of stars of M or S spectral type, excluding of course the Mno stars of Table 1. Their distribution in the parameter space is displayed in Figure 3. As can be seen in all panels of the figure, there is no clear distinction between the regions covered by S spectral types and by M spectral types. This is disappointing as it prevents using effectively the important milestone of the M to S transition as a tool to clarify the relation between the evolutions of the light curve in parameter space and of the star along the AGB. It suggests that such transition may occur at very different stages of the TP-AGB, depending on the star properties, mostly its initial mass. Table 3 lists all curves of our sample, excluding the Mno curves listed in Table 1, that have an M or S spectral type and for which the light curve has a width parameter 0.3. Their distribution in the parameter space is displayed in Figure 3. As decreases, the period, the oscillation amplitude and the colour index tend to increase. Remarkably, a dozen of curves display large irregularity parameters, well in excess of 0.6, in strong contrast with the Mno sample where a single curve, that of R Cas, has a large value of ; it has also extreme values of , and , suggesting that it is close to evolving to an Myes spectral type. The presence of a large group of irregular curves in the present sample suggests that some are evolving toward a state of unstable pulsation and might be leaving the domain of stable pulsation before becoming carbon-rich stars. Indeed, F. L. Schöier & H. Olofsson (2001), in a study of a large sample of bright carbon stars, have shown that only a quarter of them pulsate as Mira variables.
@@
We now turn to the study of the curves having 0.3 of stars of M or S spectral type, excluding of course the Mno stars of Table 1. They are listed in Table 4 and Figure 4 displays their distribution in the parameter space. At first glance, Figure 4 seems to suggest that the 0.3 curves are successors of the Mnoi-j curves having 4, while the 0.3 curves are successors of the Mnoi-j curves having 3; moreover, the location of the 0.3 curves between the Mno and carbon-rich samples may suggest that the stars having 0.3 curves will evolve to the carbon-rich variables of the present sample. Yet, several distributions, in particular in the vs and vs planes, show continuity between the Table 3 (0.3) and Table 4 (0.3) samples, raising questions on the opportunity to make a distinction between them. This is further illustrated in Section 5, where both samples are considered jointly. A close inspection of the properties of the curves listed in Table 4 is therefore necessary to understand whether such crude impressions are or not justified.
@@
The trend toward low colour indices is already apparent in the Table 4 sample of the twelve curves having 0.4. Indeed, out of the 548 curves selected by P. Merchán Benítez et al. (2023), 40 have [3.4]-[22]1.25, of which 19 are of S spectral type, 18 of M spectral type and 3 of C spectral type. Of these, 10 are part of the Table 4 sample and only 5, all of S spectral type with between 0.24 and 0.27, are part of the much larger Table 3 sample. This is evidence for a group of stars of very low colour indices, mostly of S, but also Myes spectral types, having curves of large parameter, to own a specific identity that must be associated with well-defined physical properties. A likely interpretation is offered by O. Hashimoto et al. (1998) in the case of R Hya, a star of the Table 4 sample having an M6- 9 spectral type, a colour index 0.9 and a broad curve profile, 0.36. The claim is that mass-loss has recently nearly stopped, causing the circumstellar envelope to expand away from the star and become too cold for the silicate dust grains to significantly emit in the infrared. Recent observations and analyses (L. Decin et al., 2008; A. Baudry et al., 2023; M. Joyce et al., 2024; R. Zhao-Geisler et al., 2012) have confirmed the validity of this interpretation and suggested that the cause of the absence of significant mass-loss is the likely occurrence of a thermal pulse some 200 years ago.
@@
In summary, an interpretation of the Table 4 curves as associated with large initial mass stars near the end of the TP-AGB, having light curves evolving from early Mno type (4) to the carbon-rich Table 5 sample is plausible but far from being firmly established and raises several questions that deserve careful consideration. The strong evidence for a trend toward low colour indices when the star evolves on the AGB from M to S spectral types, is particularly puzzling in this respect: it goes against an evolution toward the curves of C spectral type, which have instead larger colour indices. The argument proposed by P. Marigo et al. (2020), concerning the expected scarcity of dust when C/O approaches unity, is very relevant but raises questions concerning its detailed application, not only for the curves of Table 4 but also for those of Table 3.
@@ 3.4 Carbon-rich stars
The light curves of the carbon-rich stars of the present sample are listed in Table 5 and their distribution in the parameter space is illustrated in Figure 4. The width parameter exceeds 0.3. A clear correlation is seen between and and a weaker one between and and and . The analyses presented in the preceding subsections have left open a reliable identification of their progenitors. They found plausible that they include members of the Table 4 sample, with 0.3. But they could not exclude that they would also include members of the Table 3 sample, with 0.3. In particular, transition from Table 3 to Table 5 might occur via a regime of unstable pulsation. Both 0 and 1 curves are curves of constant luminosity. In the case of the former, the light spikes have become so narrow that they have disappeared, in spite of possibly having a large amplitude. In the case of the latter, the dips have become so narrow that they have disappeared, no matter how deep they were. A curve of constant luminosity is no longer variable and has lost memory of what it was before: if it becomes variable again, it may do it with low amplitude ( small, irrelevant) or with short spikes of light ( close to 0, irrelevant) or with short dips ( close to 1, irrelevant).
@@
J. Bergeat & L. Chevallier (2005) have collected data related to a large sample of carbon-rich stars, of which 15 are part of the Table 5 sample. Figure 5 displays the dependence of the parameters and of their light curves on effective temperature and mass loss rate . If one excludes four low curves that are close to the domain covered by the Table 4 sample (0.3), V CrB, U Cyg, RZ Peg and R Ori, tends to increase with temperature and decrease with mass loss rate while displays the opposite trend. Such a behaviour would therefore suggest that the stars having large light curves evolve in the direction of decreasing when becoming cooler. This trend is confirmed by the dependence of on the spectral index for spectral types C: for 6, namely warmer stars, 0.530.07 while for 6, namely for cooler stars, 0.390.04. This is however excluding R For, a clear outlier having a C4 spectral type and 0.35.
@@
To the extent that we can take as granted that carbon stars keep cooling down (P. Marigo, 2002), this suggests that the carbon-rich stars of the Table 5 sample having curves of large value enter it from the large side and evolve toward smaller values of . If such is the case, they would probably have a Table 3 (0.3) origin: the larger , the smaller the oscillation amplitude (Figure 4), namely the closer from a regime of unstable pulsation. Figure 6 illustrates what such a transition would look like by displaying projections on the vs , vs , vs and vs planes of the distributions in parameter space of the Table 3 and Table 5 samples. It shows that, apart from and , a transition from Table 3 to Table 5 is reasonably continuous in the other parameters. The discontinuity in amplitude is not surprising: absorption of the dust envelope is known to be larger for carbon-rich rather than oxygen-rich stars. One may instead assume that Table 3 stars such as Cyg will become irregular carbon-rich stars absent from catalogues of Mira variables. This would leave the identity of the precursors of the large members of the Table 5 curves unanswered, but one might similarly assume that they are irregular curves of S spectral type stars. In this context, one might similarly argue that the low colour index stars of the Table 4 sample, with S spectral types, would also evolve toward irregular carbon-rich stars. These comments illustrate the difficulty to reliably interpret what is observed.
@@
RS Cyg and V Oph, both parts of the Figure 5 sample, have large values of , 0.66 and 0.52, respectively. They would then be likely successors of the Table 3 sample. RS Cyg has a high temperature of 3100 K and a low mass loss rate of 210-7 M⊙yr-1 (R. Guandalini et al., 2006). Its light curve has been extensively studied and commented by R. R. Cadmus (2022, 2024). It often displays a hump on the ascending branch, close to maximum light, possibly producing a double-maximum profile. V Oph has been the target of detailed VLT observations in the infrared (K. Ohnaka et al., 2007; G. Rau et al., 2019; J. Hulberg et al., 2025) suggesting an effective temperature of 2700 K while Figure 5 quotes a temperature of 3010 K and a mass loss rate of 1.410-7 M⊙yr-1. However, the identification of the origin of curves having lower values of turns out to be difficult. BH Cru and LX Cyg were shown by S. Uttenthaler et al. (2011, 2016) to have recently switched from spectral type S to C: one would naively expect to find them near an edge of the domain covered by the carbon-rich sample in the parameter space; but such is not the case, their curves have values of (0.47 and 0.46, respectively) that sit in the middle of those of the Table 5 sample, a bit on the high side: they may be successors of the Table 3 (0.3) sample as well as of the Table 4 (0.3) sample, although the value of the temperature (3000 K, A. A. Zijlstra et al., 2004) rather favours the former.
@@
S Cep, with 0.40, has an effective temperature of 2100 K (W. Nowotny et al., 2005, 2010), leaving open both a Table 3 (0.3) and a Table 4 (0.3) origin. Such is indeed the case for all curves having values of this order or smaller. The 13 curves having 0.4 cover rather uniformly the parameter space, failing to reveal a clear separation between two groups of curves. These examples illustrate the difficulty of producing reliable interpretations of the evolution of the carbon-rich stars having curves in the Table 5 sample: the likely coexistence of Table 3 (0.3) and Table 4 (0.3) successors is a major complication and a reliable distinction between them is difficult.
@@
Finally, we recall that we have shown in Paper I (Section 5, Figure 14) the dependence of and on and remarked that carbon-rich stars and oxygen-rich stars follow similar trends, the former having slightly lower values of and of than the latter for a same value of . Figure 7 illustrates this issue in the present sample, displaying separately the samples of Tables 1, 3, 4 and 5. The Table 1 (Mno) distributions illustrate the transition from what was defined as first and second families of Mno curves in Paper I, with both and increasing; the location of the curve of R Cas (0.84) confirms the suggestion that was made in Subsection 3.2 that the star is in the process of switching to the Myes spectral type. The same trend is observed in the distributions of the Table 3 (0.3) sample, with a global shift toward larger and ; a clear shift toward large values of gives evidence for the increased instability of the pulsation regime. The concentration in small values of both and illustrates the difference between the curves of the Table 4 (0.3) sample and the others. The distributions displayed by the Table 5 sample confirm the comments made in Paper I.